OpenStudy (anonymous):

a flat rectangular piece of aluminum has a perimeter of 58 inches. the length is 7 inches longer than the width. find the width.

ganeshie8 (ganeshie8):

say length = l, width = w

ganeshie8 (ganeshie8):

do u remember the formula for 'perimeter of rectangle' ?

OpenStudy (anonymous):

P=2L+2W

ganeshie8 (ganeshie8):

yes ! 2L+2W = 58 -------------(1)

OpenStudy (anonymous):

let x+7 = length let x = width add x+x+(x+7) + (x+7) =4x +14 =?

ganeshie8 (ganeshie8):

it is given that length is 7 inches longer than the width. L = W + 7 -----------(2)

ganeshie8 (ganeshie8):

two equation, two unknowns, you can solve them

OpenStudy (anonymous):

2(x+7) + 2(x) =58 solve for x

OpenStudy (anonymous):

x=11

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Hint: x = side length

OpenStudy (anonymous):

im getting 11 for x

OpenStudy (anonymous):

11*4 = 44 44 + 14 = 58

OpenStudy (anonymous):

If you wanted to use x for length then you would represent "x-7 for the side, it can be done either way

OpenStudy (anonymous):

got it! thank you!